# Collaborative Learning in Agentic Systems: A Collective AI is Greater Than the Sum of Its Parts

### v1

@@ 2.5 Summary
The complete MOSAIC algorithm executes the processes described above in a loop. Figure 1 illustrates the algorithmic pipeline with each of the components. MOSAIC computes embeddings from batches of state-action-reward (SAR) data collected under the current policy, and these embeddings evolve over time as the policy improves. Knowledge reuse is governed by a linear combination of task-specific masks selected using similarity (computed from embeddings) and performance heuristics, with beta parameters re-initialized at each communication event and then fine-tuned to reflect downstream utility to the current task. While not part of this study, Appendix B provides a formalization of how masks can be linearly combined and how beta parameters can be initialized to extend MOSAIC to a fully-fledged lifelong reinforcement learning setting Ben-Iwhiwhu et al. (2023); Xie and Finn (2022).
Algorithm 1 provides an outline of the overall process. Pseudo-code for how the forward pass is computed is provided in Appendix E Algorithm 2. MOSAIC makes use of parallelized processes to achieve asynchronicity between learning and sharing knowledge. These parallel processes are described in Appendix E Algorithms 3, 4, and 5, which handle the updating of the mask and linear combination parameters on every communication event and manage the communication client/server, respectively. Queries and masks are shared around the system via multi-processing queues $Q_{\text{emb}}$ and $Q_{\text{mask}}$. Appendix E Algorithm 6 shows the computation of the Wasserstein task embeddings. The Python code of the complete algorithm to reproduce the results is provided at https://github.com/DMIU-ShELL/MOSAIC.
@@
7:    ReplacePeerMasks($Q_{\text{mask}}$) $\triangleright$ Receives and applies mask updates Alg. 3
8:    CommunicationModule($Q_{\text{emb}},Q_{\text{mask}}$) $\triangleright$ Handles queries and peer responses Alg. 4
@@
10:     Roll out policy $\pi_{\tau}=\pi_{\Phi\odot g(\phi_{\tau}^{\text{lc}})}$ for $z$ $\triangleright$ Forward pass Alg. 2
@@
12:     Update $\phi_{\tau}$ using policy gradient on loss $\mathcal{L}(\pi_{\tau})$triangleright Via Eq. 8
@@
15:         Sample $N$ samples from $\mathcal{D}_{\tau}$ and compute task embedding $v_{\tau}\leftarrow\psi(\mu_{\tau})$$\triangleright$ Alg. 6

### v2

@@ Appendix E Implementation Details
In this section we provide implementation details for the MOSAIC algorithm. MOSAIC computes embeddings from batches of state-action-reward (SAR) data collected under the current policy. These embeddings evolve over time as the policy improves. Knowledge reuse is governed by a linear combination of task-specific masks selected using similarity and performance heuristics. The linear combination is weighted by learned $\beta$ parameters which are re-initialized at each communication event, and then fine-tuned again to reflect downstream utility to the current task.
Algorithm 4 provides pseudocode for how MOSAIC computes the forward pass using a linear combination of masks and the binarization step. Algorithm 1 provides an outline of the main algorithmic pipeline. The agent additionally, runs two parallel threads to run and maintain interaction communication module and model while maintaining asynchronous operation. This enables MOSAIC to learn and communicate concurrently. Pseudocode for these threads are provided in Algorithms 2 and 3. The server to accept incoming communications is started as a parallel process inside the communication module, described in Algorithm 5. Queries and masks are shared around the system via multi-processing queues $Q_{\text{emb}}$ and $Q_{\text{mask}}$. Algorithm 6 shows the computation of the Wasserstein task embeddings.
System 1
OS Ubuntu 22.04.5 LTS
CPU AMD EPYC 9634 (168 logical cores)
GPU(s) 2× NVIDIA A100 80GB PCIe
GPU Driver 535.216.03
CUDA Version 12.2
RAM 512 GB
System 2
OS Ubuntu 22.04.5 LTS
CPU AMD EPYC 7713P (128 logical cores)
GPU(s) 2× NVIDIA A100 40GB PCIe
GPU Driver 535.216.03
CUDA Version 12.2
RAM 256 GB
System 3
OS Ubuntu 22.04.5 LTS
CPU AMD Ryzen 9 5950X (32 logical cores)
GPU(s) 1× NVIDIA GeForce RTX 4090
GPU Driver 535.183.01
CUDA Version 12.2
RAM 128 GB
Module Hyperparameter Value (ISL, MiniGrid, MiniHack)
PPO Learning rate 0.00025
Backbone seed 9157
Number of workers 1
Discount factor 0.99
GAE Tau 0.99
Entropy 0.01
Rollout length 512 (ISL), 2048 (MiniGrid, MiniHack)
Optimization iterations 8
Mini-batches 64
PPO ratio clip 0.1
Gradient clip 5
Total training steps 102,400 (ISL), 1,024,000 (MiniGrid), 2,048,000 (MiniHack)
Comm. Query frequency (iterations) 10 (ISL, MiniHack), 25 (MiniGrid)
Query wait (ms) 0.3
Mask wait (ms) 0.3
Embedding Reference size ($M$) 50
Number of samples ($N$) 128
Frequency (iterations) 1
Table 4: Hyperparameters used in the experiments presented in this paper. Values are provided for the reinforcement learning algorithm, the communication module and embedding module.
Network Architecture Details
Fully Connected Feature extractor: Linear (input, 200) $\rightarrow$ Linear (200, 200) $\rightarrow$ Linear (200, 200)
Critic: Linear (200, 1)
Actor: Linear (200, action)
Convolutional Feature extractor: Conv (input, 16, kernel 8, stride 4) $\rightarrow$ Conv (16, 32, kernel 4, stride 2) $\rightarrow$ Linear (8192, 256)
Critic: Linear (256, 1)
Actor: Linear (256, action)
Table 5: Network architectures used in experiments. Image sequence learning and MiniGrid benchmarks were run using the full connected network. MiniHack experiments were run using the convolutional network
Package/Framework Version
Python 3.9.13
PyTorch 2.4.0
TorchVision 0.19.0
TorchAudio 2.4.0
NumPy 1.24.3
SciPy 1.10.1
Pandas 2.2.2
Scikit-learn 1.5.1
Gym 0.26.0
Gym-Minigrid 1.2.2
Gym-CTGraph 0.1 (dev)
OpenCV-Python 4.10.0.84
Matplotlib 3.9.2
Seaborn 0.13.2
TensorBoardX 2.6.2.2
POT (Optimal Transport) 0.9.4
CUDA Toolkit 12.1.105
cuDNN 9.1.0
Table 6: Software packages and frameworks used for reproducibility.
@@
7:ReplacePeerMasks($Q_{\text{mask}}$) $\triangleright$ Receives and applies mask updates Alg. 2
8:CommunicationModule($Q_{\text{emb}},Q_{\text{mask}}$) $\triangleright$ Handles queries and peer responses Alg. 3
@@
[…]
@@
12:  Update $\phi_{\tau}$ using policy gradient on loss $\mathcal{L}(\pi_{\tau})$ $\triangleright$ Via Eq. 7
@@
15:   Sample NN samples from $\mathcal{D}_{\tau}$ and compute task embedding $v_{\tau}\leftarrow\psi(\mu_{\tau})\triangleright$ Alg. 6